# Bozen-Bolzano, IT SkillFormer

### v2

@@ 5.1 Overall Accuracy and Efficiency
Table 2: Training hyperparameters and overall accuracy for SkillFormer across view configurations. Fixed across all runs: epochs = 4, batch size = 16, output dim = 768, attention heads = 16. Hid denotes the dimensionality of the intermediate representation used in the CrossViewFusion module. It controls the capacity of the feed-forward transformation and the gating mechanism.
@@
5.2 Efficiency and Configuration Analysis
Table 2 details the hyperparameter configurations adopted across different multi-view setups. As the number of views increases, our design prioritizes efficiency without compromising accuracy. To this end, we strategically reduce the number of frames per view—preserving the temporal span with fewer sampled tokens—while proportionally increasing the LoRA rank and hidden dimension in the CrossViewFusion module. This trade-off ensures that the model retains sufficient capacity to capture inter-view dependencies, even under tighter per-view bandwidth.
5.3 Per-Domain Accuracy Analysis
Table 3: Per-scenario accuracy (%) for the majority-class baseline, the baseline models from [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray], and SkillFormer across different view configurations. Bold indicates the best-performing method per scenario, while underlined values represent the second-best. SkillFormer consistently outperforms other models in multi-view setups, particularly in Basketball, Cooking, and Bouldering.
@@ 6 Limitations and Future Work
Fixed frame sampling may miss critical temporal cues in variable activities, a limitation that could be addressed through adaptive sampling with attention-based keyframe detection. SkillFormer also underperforms significantly in subjective domains like Dancing, suggesting challenges with stylistic variability that domain adaptation strategies such as curriculum learning could improve.

### v3

@@ 5.1 Overall Accuracy and Efficiency
Table 1 compares overall accuracy between SkillFormer and baseline models from the EgoExo4D benchmark [12]. In contrast to baselines—which train separate models for egocentric and exocentric inputs and perform late fusion at inference—SkillFormer uses a single unified model for each configuration (Ego, Exos, Ego+Exos). Furthermore, we do not apply multi-crop testing, simplifying inference and reducing computational overhead.
@@
5.2 Scaling Strategy and Parameter Analysis
Table 2 details our scaling strategy across view configurations. As the number of views increases, our design prioritizes efficiency without compromising accuracy. To this end, we strategically reduce the number of frames per view (32→16)—preserving the temporal span with fewer sampled tokens—while proportionally increasing the LoRA rank (32→64), alpha (64→128), and hidden dimension (1536→2560). This trade-off compensates for reduced per-view tokens by enabling richer cross-view transformations.
Table 2: SkillFormer scaling strategy across view configurations. As views increase, we reduce frames while proportionally increasing LoRA rank and fusion capacity to maintain model expressiveness within computational constraints. Fixed across all runs: epochs = 4, batch size = 16, output dim = 768, attention heads = 16. Hid denotes the dimensionality of the intermediate representation used in the CrossViewFusion module.
@@ 5.2 Scaling Strategy and Parameter Analysis
Table 3: Per-scenario accuracy (%) for the majority-class baseline, the baseline models [12], and SkillFormer across different view configurations. Bold indicates the best-performing method per scenario, while underlined values represent the second-best. SkillFormer consistently outperforms other models in multi-view setups, particularly in Basketball, Cooking, and Bouldering.
@@ 6 Limitations and Future Work
Fixed frame sampling may miss critical temporal cues in variable activities, a limitation that could be addressed through adaptive sampling with attention-based keyframe detection or sport-tailored sampling methods [16]. SkillFormer also underperforms significantly in subjective domains like Dancing, suggesting challenges with stylistic variability that domain adaptation strategies such as curriculum learning could improve.